### v1

@@ Abstract
    We design an $O(2^{k}n^{2k}poly(n))$-time algorithm to find a maximum clique in a $n$-vertex disk graph with $k$ different sizes of radii. This is polynomial for every fixed $k$, and thus settles the open question for the case when $k=2$.
@@
    We give an $O(2^{k}n^{2rk}poly(n,r))$-time algorithm to find a maximum clique in a $n$-vertex ball graph with $k$ different sizes of radii where the ball centers lie on $r$ parallel planes. This is polynomial for every fixed $k$ and $r$, and thus contrasts the previously known NP-hardness result for finding a maximum clique in an arbitrary ball graph.
@@ 1 Introduction
The maximum clique problem is NP-hard for many well-known intersection graph classes such as intersection graph of rays [11], grounded strings [29], triangles [2], ellipses [2], and a combination of axis-aligned rectangles and unit disks [6, 18], whereas polynomial-time solvable for circle graphs [32], trapezoid graphs [20], circle trapezoid graphs [20], unit disk graphs [14], axis-aligned rectangle intersection graphs [26], and so on. However, determining the time complexity of computing a maximum clique in a disk graph is a long-standing open question [2, 3, 21]. In a seminal paper published in 1990, Clark, Colbourn, and Johnson [14] gave a polynomial-time algorithm for the case of unit disk graphs. However, the time-complexity question for general disk graphs has remained open since then. Although this was not posed as an open problem in [14], it has long been known to be a challenging problem [3], even before it was explicitly posed as an open problem (e.g., by Fishkin [21], Ambühl and Wagner [2] and Cabello [9, 10]). In fact, the problem of extending the unit-disk results in a more general setting has been emphasized in various ways in the literature, e.g., as “an intriguing open question”, as “a notorious open question in computational geometry” [4], as being “elusive with no new positive or negative results” [3], and several times as a “long-standing open problem” [18, 22, 23].
@@
Disk graph with $k$ Different Sizes of Radii: We give an $O(2^{k}n^{2k}(f(n)+n^{2}))$-time algorithm to find a maximum clique in a disk graph where $k$ is the number of different types of radii and $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph. Since $f(n)$ is polynomial in $n$ [24], for every fixed $k$, the running becomes polynomial. This settles the open question posed by Cabello [9, 10] on whether a polynomial-time algorithm exists when $k=2$.
@@
Ball graphs with $k$ Different Sizes of Radii: Given a set of $n$ balls in the Euclidean plane with $k$ different sizes of radii where the ball centers lie on $r$ parallel planes, we show how to compute a maximum clique in the corresponding ball graph in $O(2^{k}n^{2rk}(f(n)+n^{2}r))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph. For fixed $k$ and $r$, the running time becomes polynomial in $n$. We then show that the restriction of the planes being parallel can be removed as long as the input planes are all perpendicular to a different plane. This result contrasts the known NP-hardness result [4] for finding a maximum clique in a ball graph where the radii of the balls are very close to each other, i.e., in the interval $[1,1+\varepsilon]$, where $\varepsilon>0$.
@@ 3 Maximum Clique in Disk Graphs with $k$ Different Radii
In this section we give an $O(2^{k}n^{2k}(f(n)+n^{2}))$-time algorithm to find a maximum clique in a disk graph where $k$ is the number of different types of radii and $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
@@
We now consider the running time. There are at most (${n\choose 1}+{n\choose 2}=O(n^{2})$ ways to guess the pair of disks for a particular disk type, irrespective of whether these two disks are distinct or not. Therefore, the number of ways we can guess $k$ pairs is $O(n^{2k})$. For each guess of $k$ pairs, it is straightforward to construct the sets $X$ and $Y$ in $O(kn)\in O(n^{2})$ time, and the disk graph determined by the disks $(X\cup Y)$ in $O(n^{2})$ time. Let $f(n)$ be the time for computing a maximum matching in a $n$-vertex bipartite graph. Then the running time is $O(n^{2k}(f(n)+n^{2}))$. After taking the $2^{k}$ different guesses for radii sizes that may appear in a maximum clique into consideration, the overall running time becomes $O(2^{k}n^{2k}(f(n)+n^{2}))$.
@@
Given a set of $n$ disks in the Euclidean plane with $k$ different types of radii, a maximum clique in the corresponding disk graph can be computed in $O(2^{k}n^{2k}(f(n)+n^{2}))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
Note that this can easily be turned into an $O(n^{t}\cdot n^{2t}(f(n)+n^{2}))$-time algorithm to find a maximum clique with at most $t$ different types of radii in a general disk graph, by checking for all $O(n^{t})$ options for choosing $t$ different radii from $n$ possible options.
Corollary 3.2.1
Given a set of $n$ disks in the Euclidean plane with arbitrary radii and a positive integer $t$, a maximum clique with at most $t$ different radii can be computed in $O(n^{3t}(f(n)+n^{2}))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
@@ 5 Maximum Clique in Ball Graphs with $k$ Different Radii
We now give an $O(2^{k}n^{2rk}(f(n)+n^{2}r))$-time algorithm for computing a maximum clique in a ball graph, where the centers of the balls are contained in $r$ planes which are all perpendicular to a different plane (Section 5.2). For simplicity, we first consider a scenario when the input planes are parallel to each other (Section 5.1).
@@ 5.1 Ball Centers are on $r$ Parallel Planes
We now consider the running time. For a single plane, there are at most $O(n^{2})$ ways to guess the pair of balls for a particular ball type, and thus $O(n^{2k})$ ways for $k$ pairs considering all types. Since we have $r$ planes, the total number of guesses is $O(n^{2rk})$. For each guess of $rk$ pairs, it is straightforward to construct the sets $X$ and $Y$ in $O(rkn)\in O(n^{2}r)$ time, and the ball graph determined by the balls $(X\cup Y)$ in $O(n^{2})$ time. Let $f(n)$ be the time for computing a maximum matching in a $n$-vertex bipartite graph. Then the running time is $O(n^{2rk}(f(n)+n^{2}r))$. After taking the $2^{k}$ different guesses for radii sizes that may appear in a maximum clique into consideration, the overall running time becomes $O(2^{k}n^{2rk}(f(n)+n^{2}r))$.
@@
Given a set of $n$ balls in the Euclidean plane with $k$ different types of radii where the ball centers are contained on $r$ parallel planes, a maximum clique in the corresponding ball graph can be computed in $O(2^{k}n^{2rk}(f(n)+n^{2}r))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
@@ 5.2 Ball Centers are on $r$ Planes which are Perpendicular to a Different Plane
Given a set of $n$ balls in the Euclidean plane with $k$ different types of radii where the ball centers are contained on $r$ planes that are perpendicular to a single different plane, a maximum clique in the corresponding ball graph can be computed in $O(2^{k}n^{2rk}(f(n)+n^{2}r))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
@@ 6.2 Direction for Future Research
In this paper we gave an $O(2^{k}n^{2k}poly(n))$-time algorithm to find a maximum clique in a $n$-vertex disk graph with $k$ different radii, and an $O(2^{k}n^{2rk}poly(n,r))$-time algorithm to find a maximum clique in a $n$-vertex ball graph with $k$ different radii where the ball centers are contained in $r$ planes that are perpendicular to a single different plane. Designing faster algorithms would be a natural direction to explore. One may attempt to remove the perpendicularity constraint for the case of ball graphs. The lower bound on the time complexity is another avenue to explore. The NP-hardness reduction for computing a maximum clique in a ball graph [5] allows for arbitrary radii whereas our result provides polynomial-time algorithms when $r\in O(1)$ and $k\in O(1)$. Therefore, it would be interesting to examine cases when $k\in o(n)$.

### v2

@@ Abstract
    We design an $O(n^{2k}poly(n))$-time algorithm to find a maximum clique in a $n$-vertex disk graph with $k$ different sizes of radii. This is polynomial for every fixed $k$, and thus settles the open question for the case when $k=2$.
@@
    We give an $O(n^{2rk}poly(n,r))$-time algorithm to find a maximum clique in a $n$-vertex ball graph with $k$ different sizes of radii where the ball centers lie on $r$ parallel planes. This is polynomial for every fixed $k$ and $r$, and thus contrasts the previously known NP-hardness result for finding a maximum clique in an arbitrary ball graph.
@@ 1 Introduction
The maximum clique problem is NP-hard for many well-known intersection graph classes such as intersection graph of rays [11], grounded strings [29], triangles [2], ellipses [2], and a combination of axis-aligned rectangles and unit disks [6, 18], whereas polynomial-time solvable for circle graphs [32], trapezoid graphs [20], circle trapezoid graphs [20], unit disk graphs [14], axis-aligned rectangle intersection graphs [26], and so on. However, determining the time complexity of computing a maximum clique in a disk graph is a long-standing open question [2, 3, 21]. In a seminal paper published in 1990, Clark, Colbourn, and Johnson [14] gave a polynomial-time algorithm for the case of unit disk graphs. However, the time-complexity question for general disk graphs has remained open since then. Although this was not posed as an open problem in [14], it has long been known to be a challenging problem [3], even before it was explicitly posed as an open problem (e.g., by Fishkin [21], Ambühl and Wagner [2] and Cabello [9, 10]). In fact, the problem has been emphasized in various ways in the literature, e.g., as “an intriguing open question”, as “a notorious open question in computational geometry” [4], as being “elusive with no new positive or negative results” [3], and several times as a “long-standing open problem” [18, 22, 23].
@@
Disk graph with $k$ Different Sizes of Radii: We give an $O(n^{2k}(f(n)+n^{2}))$-time algorithm to find a maximum clique in a disk graph where $k$ is the number of different types of radii and $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph. Since $f(n)$ is polynomial in $n$ [24], for every fixed $k$, the running becomes polynomial. This settles the open question posed by Cabello [9, 10] on whether a polynomial-time algorithm exists when $k=2$.
@@
Ball graphs with $k$ Different Sizes of Radii: Given a set of $n$ balls in the Euclidean plane with $k$ different sizes of radii where the ball centers lie on $r$ parallel planes, we show how to compute a maximum clique in the corresponding ball graph in $O(n^{2rk}(f(n)+n^{2}r))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph. For fixed $k$ and $r$, the running time becomes polynomial in $n$. We then show that the restriction of the planes being parallel can be removed as long as the input planes are all perpendicular to a different plane. This result contrasts the known NP-hardness result [4] for finding a maximum clique in a ball graph where the radii of the balls are very close to each other, i.e., in the interval $[1,1+\varepsilon]$, where $\varepsilon>0$.
@@ 3 Maximum Clique in Disk Graphs with $k$ Different Radii
In this section we give an $O(n^{2k}(f(n)+n^{2}))$-time algorithm to find a maximum clique in a disk graph where $k$ is the number of different types of radii and $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
@@
We now consider the running time. There are at most (${n\choose 0}+{n\choose 1}+{n\choose 2}=O(n^{2})$ ways to guess the pair of disks for a particular disk type, irrespective of whether these two disks are distinct or not. Therefore, the number of ways we can guess $k$ pairs is $O(n^{2k})$. For each guess of $k$ pairs, it is straightforward to construct the sets $X$ and $Y$ in $O(kn)\in O(n^{2})$ time, and the disk graph determined by the disks $(X\cup Y)$ in $O(n^{2})$ time. Let $f(n)$ be the time for computing a maximum matching in a $n$-vertex bipartite graph. Then the running time is $O(n^{2k}(f(n)+n^{2}))$.
@@
Given a set of $n$ disks in the Euclidean plane with $k$ different types of radii, a maximum clique in the corresponding disk graph can be computed in $O(n^{2k}(f(n)+n^{2}))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
@@ 5 Maximum Clique in Ball Graphs with $k$ Different Radii
We now give an $O(n^{2rk}(f(n)+n^{2}r))$-time algorithm for computing a maximum clique in a ball graph, where the centers of the balls are contained in $r$ planes which are all perpendicular to a different plane (Section 5.2). For simplicity, we first consider a scenario when the input planes are parallel to each other (Section 5.1).
@@ 5.1 Ball Centers are on $r$ Parallel Planes
We now consider the running time. For a single plane, there are at most $O(n^{2})$ ways to guess the pair of balls for a particular ball type, and thus $O(n^{2k})$ ways for $k$ pairs considering all types. Since we have $r$ planes, the total number of guesses is $O(n^{2rk})$. For each guess of $rk$ pairs, it is straightforward to construct the sets $X$ and $Y$ in $O(rkn)\in O(n^{2}r)$ time, and the ball graph determined by the balls $(X\cup Y)$ in $O(n^{2})$ time. Let $f(n)$ be the time for computing a maximum matching in a $n$-vertex bipartite graph. Then the running time is $O(n^{2rk}(f(n)+n^{2}r))$.
@@
Given a set of $n$ balls in the Euclidean plane with $k$ different types of radii where the ball centers are contained on $r$ parallel planes, a maximum clique in the corresponding ball graph can be computed in $O(n^{2rk}(f(n)+n^{2}r))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
@@ 5.2 Ball Centers are on $r$ Planes which are Perpendicular to a Different Plane
Given a set of $n$ balls in the Euclidean plane with $k$ different types of radii where the ball centers are contained on $r$ planes that are perpendicular to a single different plane, a maximum clique in the corresponding ball graph can be computed in $O(n^{2rk}(f(n)+n^{2}r))$ time. Here $f(n)$ is the time to compute a maximum matching in a $n$-vertex bipartite graph.
@@ 6.2 Direction for Future Research
In this paper we gave an $O(n^{2k}poly(n))$-time algorithm to find a maximum clique in a $n$-vertex disk graph with $k$ different radii, and an $O(n^{2rk}poly(n,r))$-time algorithm to find a maximum clique in a $n$-vertex ball graph with $k$ different radii where the ball centers are contained in $r$ planes that are perpendicular to a single different plane. Designing faster algorithms would be a natural direction to explore. One may attempt to remove the perpendicularity constraint for the case of ball graphs. The lower bound on the time complexity is another avenue to explore. The NP-hardness reduction for computing a maximum clique in a ball graph [5] allows for arbitrary radii whereas our result provides polynomial-time algorithms when $r\in O(1)$ and $k\in O(1)$. Therefore, it would be interesting to examine cases when $k\in o(n)$.
@@
Acknowledgements
The work is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC). We thank Soichiro Yamazaki for useful feedback.